Please help!!! Thanks in advance for you time!
A three-stage rocket has the following probabilities for failure: The probability for failure at stage one is 1/6; at stage two,1/8; and at stage three,1/10. What is the probability of a successful flight, given that the first stage was successful?

Answer by stanbon(75887)

Thanks in advance for you time!
A three-stage rocket has the following probabilities for failure: The probability for failure at stage one is 1/6; at stage two,1/8; and at stage three,1/10. What is the probability of a successful flight, given that the first stage was successful?
--------------------------------------
P(success|1st stage success) = P(all succeed)/P(1st stage succeed)
= [(5/6)*(7/8)*(9/10)]/[5/6] = (7/8)(9/10 = 63/80
============================================
Cheers,
Stan H.